# On degenerate metric resolvent: local monotonicity, restricted maximality and beyond

### v2

@@ Abstract
We study the basic properties of resolvent equipped with degenerate metric, under local monotonicity and restricted maximality, and further propose several conditions for the well-defineness and convergence of the fixed point iterations. The results help to understand the behaviours of many operator splitting algorithms in the kernel space of degenerate metric.
@@
Resolvent, degenerate metric, proximal point algorithm (PPA), operator splitting algorithms, local monotonicity, restricted maximality.
@@ 1.1 Motivation
where $\mathcal{H}$ is a real Hilbert space and $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is a set-valued operator. Built on the classical proximal point algorithm (PPA) in 1970s [25, 28], many splitting methods have been developed in the literature during the past five decades. Very recently, a degenerate preconditioned version of PPA was studied in [11], which provided a powerful framework for analyzing various existing splitting methods and also for devising new and flexible schemes. The degenerate preconditioned PPA can be described by the following fixed-point iterate
@@
where $\mathcal{Q}:\mathcal{H}\mapsto\mathcal{H}$ is a linear, bounded, self-ajoint and positive semi-definite operator. To ensure the iterate to be well-defined, the operate $\mathcal{T}$ has to be single-valued and have full domain. However, this key condition was only stated as a basic assumption in [11, Definition 2.1] and hence motivates us to find sufficient guarantees for the operator $\mathcal{T}$.
@@
Here, we adopt a different (but essentially equivalent) scheme (3) from [11, Eq.(1.6)] for two reasons: (i) (3) explicitly preserves both solution trajectory $\{u^{k}\}_{k\in\mathbb{N}}$ and shadow sequence $\{z^{k}\}_{k\in\mathbb{N}}$; (ii) the degenerate metric $\mathcal{Q}$ in (4) is easier to analyze than that one in [11, Eq.(1.5)]. Then, one may ask: Can we conclude from (4) only (without ad hoc analysis) the well-definedness of the iterations and the convergence of $\{u^{k}\}_{k\in\mathbb{N}}$? More broadly, this further motivates us to study a general setting of (2) and find the sufficient conditions on $\mathcal{A}$ and $\mathcal{Q}$, to guarantee the well-definedness of $\mathcal{T}$ and the convergence of $\{x^{k}\}_{k\in\mathbb{N}}$.
@@ 1.2 Related works
The metric resolvent $\mathcal{T}$ was also coined the term warped resolvent in [14, Definition 1.1] or $F$-resolvent [8, Definition 4.1] in various contexts. [8, Sect. 3] assumed $\mathcal{Q}$ to be maximally monotone, strictly monotone, 3 monotone and surjective; [14, Definition 1.1] assumed $\mathsf{ran}\mathcal{Q}\subseteq\mathsf{ran}(\mathcal{A}+\mathcal{Q})$ and disjoint injectivity of $\mathcal{A}+\mathcal{Q}$, and [14, Proposition 3.9] further listed specific requirements on $\mathcal{A}$ and $\mathcal{Q}$, such as surjectivity, uniform/strict/strong monotonicity. However, they are much stronger than, and fundamentally different from our degenerate setting.
To our knowledge, [17, Theorem 3.3] first studied the convergence for the limiting case of Chambolle–Pock algorithm [15], which actually corresponds to a degenerate metric $\mathcal{Q}$ (see [35, Eq.(32)] for details). This technique was later adopted in [22, Theorem 3.4], which extended the discussion to more general problems in a finite-dimensional setting. At almost the same time of [22], [12, Sect. 3] also discussed the degenerate metric for a class of preconditioned ADMM in an infinite dimensional space. [11, Sect. 2], where the authors performed a systematic study for this issue. The works of [35, 37] further exemplified the degeneracy with many practical examples and developed the reduced forms of resolvent to remove the redundancy. more theoretical results of [11] on some special cases have been reported in [7]. The degeneracy has been further extended to derive various new algorithms [36, 1, 31, 24].
@@ 1.3 Contributions
This work aims to provide a fundamental analysis and understanding of the degenerate metric resolvent. The contributions are in order.
(1) Sect. 2 studies the single-valued, $\mathcal{Q}$-firmly nonexpansive and $\mathcal{Q}$-demiclosed properties of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ under local monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$, which greatly weakens the usual monotone property of $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ in literature. Moreover, Lemma 2.14 develops several equivalent forms of $\mathcal{T}$ based on a conjugate view of $\mathcal{A}^{-1}$ via generalized Moreau’s identity.
(2) We in Sect. 3 give the full domain condition based on a conjugate view and develop a restricted Minty’s theorem based on a notion of restricted maximaly monotonicity and propose specific conditions on $\mathcal{A}$ and $\mathcal{Q}$ for the full domain. Sect. 4 further presents the sufficient and necessary conditions for the single-valuedness of $\mathcal{T}$ from both primal and conjugate perspectives. Our results extends the existing ones, from injectivity to restricted injectivity.
(3) We in Theorem 5.4 first show the weak convergence in the range space of $\mathcal{Q}$ under restricted maximal monotonicity only, and further show in Theorem 5.11 the whole convergence of $\{x^{k}\}_{k\in\mathbb{N}}$ under additional conditions. The comparisons with existing works are also discussed in details. Finally, Sect. 6 exemplifies the results with several splitting algorithms.
@@ 1.4 Notations, assumptions and preliminaries
We use standard notations and concepts from convex analysis and variational analysis, which can all be found in the classical and recent monographs [29, 30, 5, 9].
@@
As mentioned in Sect. 1.1, we focus on linear metric. More specifically, we put the following assumption on $\mathcal{Q}$ throughout the paper:
@@
(i) In particular, we assume the closedness of range of $\mathcal{Q}$, since (1) it is often satisfied by many splitting algorithms; (2) it considerably simplifies the analysis. A simple criterion for the closedness is: $\exists\alpha>0$, such that $\|x\|_{\mathcal{Q}}^{2}\geq\alpha\|x\|^{2},\ \forall x\in\mathsf{ran}\mathcal{Q}$ [5, Fact 2.26].
(ii) Assumption 1 is much weaker than the rigorous conditions of $\mathcal{Q}$ listed in [14, Proposition 3.9], e.g., strict monotonicity, surjectivity. We do not put assumption on $\mathcal{A}$ here, since the monotonicity of $\mathcal{A}$ should be closely linked to the degeneracy of $\mathcal{Q}$, and deserves particular investigation in the main part.
Before our exposition, some basic but important properties of $\mathcal{Q}$ are collected below.
@@
(i) [5, Fact 2.25] An orthogonal decomposition of $\mathcal{H}$ is: $\mathcal{H}=\mathsf{ran}\mathcal{Q}+\mathsf{ker}\mathcal{Q}$.
(ii) [5, Proposition 3.30] Any point $x\in\mathcal{H}$ can be decomposed as $x=x_{\mathsf{r}}+x_{\mathsf{k}}$, where $x_{\mathsf{r}}:=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x$ and $x_{\mathsf{k}}:=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}x$ are the $\mathsf{ran}\mathcal{Q}$- and $\mathsf{ker}\mathcal{Q}$-parts of $x$.
(iii) [10, Theorem 4] $\mathcal{Q}$ admits an (actually unique) self-adjoint square-root, denoted as $\sqrt{\mathcal{Q}}$, such that $\mathcal{Q}=\sqrt{\mathcal{Q}}\sqrt{\mathcal{Q}}$.
[…]
Fact 2
Given $\mathcal{Q}$ satisfying Assumption 1, then $\langle\cdot|\cdot\rangle_{\mathcal{Q}}$ is weak-strong continuous, i.e., if $x^{k}\rightharpoonup x^{\star}$ and $u^{k}\rightarrow u^{\star}$, then $\langle x^{k}|u^{k}\rangle_{\mathcal{Q}}\rightarrow\langle x^{\star}|u^{\star}\rangle_{\mathcal{Q}}$.
Proof 1.1.
We deduce that
$\langle x^{k}|u^{k}\rangle_{\mathcal{Q}}-\langle x^{\star}|u^{\star}\rangle_{\mathcal{Q}}$ $=$ $\langle x^{k}-x^{\star}|u^{\star}\rangle_{\mathcal{Q}}+\langle x^{k}|u^{k}-u^{\star}\rangle_{\mathcal{Q}}$
$\leq$ $\langle x^{k}-x^{\star}|\mathcal{Q}u^{\star}\rangle+\big{\|}\mathcal{Q}x^{k}\big{\|}\cdot\big{\|}u^{k}-u^{\star}\big{\|}$
$\rightarrow$ 0,since xkx0 and $0,\qquad\text{since $x^{k}-x^{\star}\rightharpoonup 0$ and $u^{k}-u^{\star}\rightarrow 0$}$
which shows that $\langle x^{k}|u^{k}\rangle_{\mathcal{Q}}-\langle x^{\star}|u^{\star}\rangle_{\mathcal{Q}}\rightarrow 0$.
2 Local monotonicity related to degenerate metric
Almost all of the existing results of metric resolvent were developed based on the usual assumption that $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is maximally monotone. This part will show that this assumption can be weakened to local monotonicity under degenerate metric.
Sect. 24 focus on the single step of (2), which is rewritten for simplicity as:
$0\in\mathcal{A}y+\mathcal{Q}(y-x)\Longleftrightarrow y\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x:=\mathcal{T}x.$ (5)
Here, we use $y\in\mathcal{T}x$, since $\mathcal{T}$ is not guaranteed to be single-valued for the moment.
2.1 Motivations and definitions
Observing (5), the monotonicity of $\mathcal{A}$ should be associated with the degeneracy of $\mathcal{Q}$, which gives rise to the following important definition, which also appeared in [11].
Definition 2.1.
@@ 2.1 Motivations and definitions
$\langle x-y|u-v\rangle\geq 0,\ \forall(x,u)\in\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q}),\ \forall(y,v)\in\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q}).$
This is similar to the notion of so-called local monotonicity [5, Definition 25.7, Proposition 25.8], which defines the monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times C)$, where $C$ is an open convex subset of $\mathcal{H}$, such that $C\cap\mathsf{ran}\mathcal{A}\neq\varnothing$. Our Definition 2.1 uses a subspace $\mathsf{ran}\mathcal{Q}$ instead of $C$. The essential difference here is that $\mathsf{int}\ \mathsf{ran}\mathcal{Q}=\varnothing$ but $\mathsf{int}\ C\neq\varnothing$. We will see in Sect. 3 that due to this difference, it is difficult to incorporate maximality into this local monotonicity.
Definition 2.1 requires $\mathcal{A}$ to be monotone in $\mathsf{ran}\mathcal{A}\cap\mathsf{ran}\mathcal{Q}$ only, regardless of the behaviours of $\mathcal{A}$ in $\mathsf{ran}\mathcal{A}\cap\mathsf{ker}\mathcal{Q}$. Clearly, this definition is much weaker than the usual (globally) monotone assumption of $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$, since nothing is enforced for $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ker}\mathcal{Q})$ in Definition 2.1.
Definition 2.1 was used in [11, Definition 2.4, Proposition 2.5] to describe $\mathcal{Q}$-monotonicity based on an inverse image of degenerate metric $\mathcal{Q}$. As a comparison, we now take a conjugate [^1] Let us explain why it is called ‘conjugate’. If $\mathcal{A}=\partial f$ for some $f\in\Gamma(\mathcal{H})$, then $\mathcal{A}^{-1}=\partial f^{*}$, where $f^{*}$ is a conjugate function of $f$. Hence, we call the analysis based on $\mathcal{A}^{-1}$ as a conjugate view. route based on $\mathcal{A}^{-1}$ and obtain the following basic observation.
Fact 1.
@@
Fact 1 transforms the local monotonicity to that of $\mathcal{A}^{-1}$. This will lead us to a conjugate perspective in later developments. This fact is an extension of [5, Proposition 20.10] to the local monotonicity within a subspace. The basic condition of $\mathsf{ran}\mathcal{A}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing$ in Definition 2.1 is also closely related to the monotonicity of $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$. Indeed, if $\mathsf{ran}\mathcal{A}\cap\mathsf{ran}\mathcal{Q}\neq\varnothing$, then $\exists x\in\mathcal{H}$, such that $\mathcal{Q}x\in\mathsf{ran}\mathcal{A}=\mathsf{dom}\mathcal{A}^{-1}$. It implies $\mathsf{dom}(\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}})\supseteq\{\mathcal{Q}x\}\neq\varnothing$. This is a very basic prerequisite for any viable operator.
2.2 Basic properties of $\mathcal{T}$ under local monotonicity
If we drop the usual assumption of $\mathcal{A}$ as being maximal monotone from $\mathcal{H}$ to $2^{\mathcal{H}}$, it should be careful to enforce new conditions on $\mathcal{A}$ as a new start, such that the requirements on $\mathcal{A}$ are kept minimal. In addition, since we has not dealt with the full domain issue of $\mathcal{T}$ (which will be discussed in Sect. 3), we use $\mathsf{dom}\mathcal{T}$ in this section. We for the moment put Definition 2.1 as
@@ 2.2 Basic properties of $\mathcal{T}$ under local monotonicity
(ii) [Local monotonicity] $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ is monotone.
We are going to show that many nice properties of $\mathcal{T}$ presented in literature still hold under local monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ only.
Many results rely on the following basic but important observation.
Fact 2.
@@
2.2.1 Single-valuedness of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$
The following simple but crucial fact will be used repeatedly in our later developments.
@@ 2.2.1 Single-valuedness of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$
Under Assumptions 1 and 2, $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ (or equivalently saying, $\mathcal{Q}\circ\mathcal{T}$) is always single-valued.
Proof 2.3.
Suppose $y_{1}$ and $y_{2}$ are two distinct outcomes of the operator $\mathcal{T}$ under the same input $x\in\mathsf{dom}\mathcal{T}$, i.e., $y_{1}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$ and $y_{2}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$. Then, we have
@@
Then,
[…]
$=$ $\big{\langle}y_{1}-y_{2}|\mathcal{Q}(x-y_{1})-\mathcal{Q}(x-y_{2})\big{\rangle}$
[…]
@@
2.2.2 $\mathcal{Q}$-firm nonexpansiveness of $\mathcal{T}$
[8, Proposition 4.2] and [12, Lemma 3.2] showed that $\mathcal{T}$ is $\mathcal{Q}$-firmly nonexpansive under maximally monotone $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ and non-degenerate $\mathcal{Q}$. This was also discussed in [34, Sect. 2]. As emphasized in [11, Lemma 2.6], this key property is preserved under local monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ for degenerate case.
Fact 3.
[11, Lemma 2.6] Under Assumptions 1 and 2, $\mathcal{T}$ given as (2) is $\mathcal{Q}$-firmly nonexpansive, i.e.,
@@ 2.2.2 $\mathcal{Q}$-firm nonexpansiveness of $\mathcal{T}$
There is still a slight difference between our result and [11, Lemma 2.6]. The latter required $\mathcal{T}$ to be well-defined, which, however, is not assumed in our result. Nevertheless, we clarify that all the expressions in above inequalities are well-defined, even if $\mathcal{T}$ is multi-valued. Indeed, $\|\mathcal{T}x\|_{\mathcal{Q}}=\|\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x\|_{\mathcal{Q}}$ and $\langle\mathcal{T}x|\mathcal{Q}y\rangle=\langle\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x|\mathcal{Q}y\rangle$, $\forall(x,y)\in\mathsf{dom}\mathcal{T}\times\mathcal{H}$. They are well-defined, since $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ is single-valued (by Lemma 2.2). To put it more explicitly, combining with Fact 2, one can equivalently rewrite (7) as $\forall x_{\mathsf{r}},y_{\mathsf{r}}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{dom}\mathcal{T})$:
@@
2.2.3 $\mathcal{Q}$-demiclosedness of $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}$
We first extend the notion of demiclosedness [5, Definition 4.26] and Browder’s demiclosedness principle [5, Theorem 4.27] to any $\mathcal{Q}$-firmly nonexpansive operator $\mathcal{S}$.

### v3

# On degenerate metric resolvent: well-definedness and new convergence analysis

@@ Abstract
We study the basic properties of resolvent equipped with degenerate metric based on restricted maximal monotonicity, and further propose several conditions for the well-definedness and weak convergence of the fixed point iterations. The results help to understand the behaviours of many operator splitting algorithms in the kernel space of degenerate metric.
@@
Resolvent, degenerate metric, restricted maximal monotonicity, proximal point algorithm (PPA), operator splitting algorithms.
@@ 1.1 Motivation
where $\mathcal{H}$ is a real Hilbert space and $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ is a set-valued operator. Built on the classical proximal point algorithm (PPA) in 1970s [24, 27], many splitting methods for solving (1) have been developed in the literature during the past five decades. Very recently, a degenerate preconditioned version of PPA was studied in [10], which provided a powerful framework for analyzing various existing splitting methods and also for devising new and flexible schemes. The degenerate preconditioned PPA can be described by the following fixed-point iterate
@@
where $\mathcal{Q}:\mathcal{H}\mapsto\mathcal{H}$ is a linear, bounded, self-ajoint and positive semi-definite operator. To ensure the iterate to be well-defined, the operate $\mathcal{T}$ has to be single-valued and of full domain. However, this key condition was only stated as a basic assumption in [10, Definition 2.1] and hence motivates us to find sufficient guarantees for the well-definedness of $\mathcal{T}$.
@@
Here, we adopt a different (but essentially equivalent) scheme (3) from [10, Eq.(1.6)] for two reasons: (i) (3) explicitly preserves both solution trajectory $\{u^{k}\}_{k\in\mathbb{N}}$ and shadow sequence $\{z^{k}\}_{k\in\mathbb{N}}$; (ii) the degenerate metric $\mathcal{Q}$ in (4) is easier to analyze than that in [10, Eq.(1.5)]. Then, one may ask: Can we conclude the well-definedness of the iterations and the convergence of $\{u^{k}\}_{k\in\mathbb{N}}$ from (4) only (without ad hoc analysis as [30] did)? More broadly, this further motivates us to study a general setting of (2) and find the sufficient conditions on $\mathcal{A}$ and $\mathcal{Q}$, to guarantee the well-definedness of $\mathcal{T}$ and the convergence of $\{x^{k}\}_{k\in\mathbb{N}}$.
@@ 1.2 Related works
When $\mathcal{Q}$ is replaced by a general non-linear operator, the operator $\mathcal{T}$ in (2) was coined the term warped resolvent in [13, Definition 1.1] and was ensured to be well-defined by requiring that $\mathsf{ran}\mathcal{Q}\subseteq\mathsf{ran}(\mathcal{A}+\mathcal{Q})$ and disjoint injectivity of $\mathcal{A}+\mathcal{Q}$ (see [12, Proposition 2.1] for instance). Besides, a group of sufficient conditions for well-definedness of $\mathcal{T}$ are listed in [13, Proposition 3.9] such as uniform/strict/strong monotonicity. However, these conditions are much strong and hard to verify. Since in this study we only focus on the operator $\mathcal{T}$ with $\mathcal{Q}$ being linear, we are able to find more mild conditions on $\mathcal{A}$ and $\mathcal{Q}$ so that the operator $\mathcal{T}$ to be well-defined.
The convergence issue of degenerate version of (2) has been first tackled in [16, 22] for specific algorithms (corresponding to particular $\mathcal{A}$ and $\mathcal{Q}$) in finite dimensional setting. However, the continuity of the proximity operator—as a main assumption proposed therein—is generally not satisfied in weak topology of infinite dimensional space. For infinite dimensional case, [11] showed the weak convergence for a class of preconditioned ADMM (associated with degenerate preconditioner) based on demiclosedness principle (under additional conditions listed in [11, Lemma 3.4]), [10] further performed a systematic study for general $\mathcal{A}$ and $\mathcal{Q}$. We in this paper for the more general scheme (2) provide more new convergence results with weakened conditions than existing works.
@@ 1.3 Contributions
This work aims to provide a fundamental analysis and understanding of the degenerate metric resolvent. The paper is organized as follows. Sect. 2 presents the basic properties of $\mathcal{T}$ under the restricted monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$, which are used as preliminary results in Sect. 35 for analyzing full domain, single-valuedness and convergence, respectively. Finally, Sect. 6 exemplifies the results with several splitting algorithms. The contributions are in order.
(1) Full domain: We in Theorem 3.1 give a sufficient and necessary condition for the full domain of $\mathcal{T}$ and further in Corollary 3.7 provide a sufficient condition that is easy to verify in practice. These results are obtained by generalized Moreau’s decomposition identity (cf. Lemma A.1) and restricted Minty’s theorem (cf. Theorem 3.4), based on a notion of restricted maximal monotonicity.
(2) Single-valuedness: Sect. 4 presents sufficient (and necessary) conditions for the single-valuedness of $\mathcal{T}$. Our results weaken the conditions of existing works, from injectivity to restricted injectivity.
(3) Convergence: Theorem 5.3 shows the weak convergence in the range space of $\mathcal{Q}$ under restricted maximal monotonicity only. This result extends [22, Theorem 3.4] to infinite dimensional setting with milder condition than [22]. Notably, the operator $\mathcal{A}$ is not assumed to be maximal monotone for this result. We further show in Theorem 5.6 the whole convergence of $\{x^{k}\}_{k\in\mathbb{N}}$ under additional conditions. The comparisons with existing works are also discussed in details.
@@ 1.4 Notations, assumptions and preliminaries
Throughout the paper, we make the following assumption on $\mathcal{Q}$.
@@
A simple criterion for the closedness is: $\exists\alpha>0$, such that $\|x\|_{\mathcal{Q}}^{2}\geq\alpha\|x\|^{2},\ \forall x\in\mathsf{ran}\mathcal{Q}$ [4, Fact 2.26].
A few basic properties of $\mathcal{Q}$ are listed here.
@@
(i) [4, Fact 2.25] An orthogonal decomposition of $\mathcal{H}$ is: $\mathcal{H}=\mathsf{ran}\mathcal{Q}+\mathsf{ker}\mathcal{Q}$.
(ii) [4, Proposition 3.30] Any point $x\in\mathcal{H}$ can be decomposed as $x=x_{\mathsf{r}}+x_{\mathsf{k}}$, where $x_{\mathsf{r}}:=\mathcal{P}_{\mathsf{ran}\mathcal{Q}}x$ and $x_{\mathsf{k}}:=\mathcal{P}_{\mathsf{ker}\mathcal{Q}}x$ are the $\mathsf{ran}\mathcal{Q}$- and $\mathsf{ker}\mathcal{Q}$-parts of $x$.
(iii) [9, Theorem 4] $\mathcal{Q}$ admits an (actually unique) self-adjoint square-root, denoted as $\sqrt{\mathcal{Q}}$, such that $\mathcal{Q}=\sqrt{\mathcal{Q}}\sqrt{\mathcal{Q}}$.
2 Basic properties under restricted monotonicity
In this part, we first introduce a restricted version of monotonicity, from which follow several well-known properties such as firm nonexpansiveness and demiclosdedness in terms of $\mathcal{Q}$-norm. These preliminary results will be essential for Sect. 35 later.
2.1 Restricted monotonicity
Observing (2), the monotonicity of $\mathcal{A}$ should be associated with the degeneracy of $\mathcal{Q}$, which gives rise to the following important definition.
Definition 2.1 (restricted monotonicity).
@@ 2.1 Restricted monotonicity
$\langle x-y|u-v\rangle\geq 0,\ \forall(x,u),(y,v)\in\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q}).$
The restricted monotonicity is equivalent to the $\mathcal{M}$-monotonicity [10, Proposition 2.5] and also closely related to the local monotonicity [4, Definition 25.7], which requires $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times C)$ to be monotone with the set $C$ being open. The essential difference between the restricted and local versions of monotonicity lies in that $\mathsf{int}\ \mathsf{ran}\mathcal{Q}$ could be empty while $\mathsf{int}\ C$ is always nonempty.
It is well-known that $\mathcal{A}$ is monotone if and only if $\mathcal{A}^{-1}$ is monotone [4, Proposition 20.22]. This fact can be extended directly as below and will provide us with a new perspective for studying the well-definedness of $\mathcal{T}$.
Proposition 2.2.
@@
For convenience of later references, we put Definition 2.1 as
@@
(ii) [Restricted monotonicity] $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ is monotone.
2.2 Basic properties
First, it is easy to recognize
Fact 1.
@@ 2.2 Basic properties
Now, we show several basic properties on $\mathcal{T}$ based on Assumption 2 and Fact 1. The first is the single-valuedness of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$.
@@
Under Assumptions 1 and 2, $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ is single-valued.
Proof 2.4.
Let $x\in\mathsf{dom}\mathcal{T}$ and assume that $y_{1}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$ and $y_{2}\in(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}x$. Then, we have
@@
Thus, by Assumptions 1 and 2, we derive that
00 0 $\geq$ $-\|y_{1}-y_{2}\|_{\mathcal{Q}}^{2}=\big{\langle}y_{1}-y_{2}|\mathcal{Q}(y_{2}-y_{1})\big{\rangle}\quad$
$=$ $\big{\langle}y_{1}-y_{2}|\mathcal{Q}(x-y_{1})-\mathcal{Q}(x-y_{2})\big{\rangle}\geq 0,$
@@
[7, Proposition 4.2] and [11, Lemma 3.2] showed that $\mathcal{T}$ is $\mathcal{Q}$-firmly nonexpansive under maximally monotone $\mathcal{A}:\mathcal{H}\mapsto 2^{\mathcal{H}}$ and non-degenerate $\mathcal{Q}$. This was also discussed in [32, Sect. 2]. As emphasized in [10, Lemma 2.6], this key property is preserved under restricted monotonicity of $\mathsf{gra}\mathcal{A}\cap(\mathcal{H}\times\mathsf{ran}\mathcal{Q})$ for degenerate case.
Fact 2.
[10, Lemma 2.6] Under Assumptions 1 and 2, $\mathcal{T}$ given as (2) is $\mathcal{Q}$-firmly nonexpansive, i.e.,
@@
There is still a slight difference between our result and [10, Lemma 2.6]. The latter required $\mathcal{T}$ to be well-defined, which, however, is not assumed in our result. Nevertheless, we clarify that all the expressions in above inequalities are well-defined, even if $\mathcal{T}$ is multi-valued. Indeed, $\|\mathcal{T}x\|_{\mathcal{Q}}=\|\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x\|_{\mathcal{Q}}$ and $\langle\mathcal{T}x|\mathcal{Q}y\rangle=\langle\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{T}x|\mathcal{Q}y\rangle$, $\forall(x,y)\in\mathsf{dom}\mathcal{T}\times\mathcal{H}$. They are well-defined, since $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}$ is single-valued (by Lemma 2.3). To put it more explicitly, combining with Fact 1, one can equivalently rewrite (6) as $\forall x_{\mathsf{r}},y_{\mathsf{r}}\in\mathcal{P}_{\mathsf{ran}\mathcal{Q}}(\mathsf{dom}\mathcal{T})$:
@@
The concept of demiclosedness [4, Definition 4.26] and Browder’s demiclosedness principle [4, Theorem 4.27] are crucial for convergence analysis, see [11, Lemma 3.4, Theorem 3.1] for example. Here it is natural (or evenly, trivial) to extend the results to any $\mathcal{Q}$-firmly nonexpansive operator $\mathcal{S}$.